# JingFang: A Traditional Chinese Medicine Large Language Model of Expert-Level Medical Diagnosis and Syndrome Differentiation-Based Treatment

### v1

@@ 4 Experiments
In this section, we explore JF’s advantages and limitations in practical applications via a series of systematic experiments. We demonstrate the performance of JF in different tasks from two core aspects: the accuracy of TCM syndrome differentiation and the effectiveness of multi-round TCM consultation. A detailed explanation of the evaluation methods and results is also provided.
4.1 Baseline Model
To comprehensively evaluate the performance of JF, we selected several representative open-source TCM models, i.e., Zhongjing-7B(Yang et al., 2024b), Mingyi-7B(Liao et al., 2024), Shenlong-7B(Wei Zhu & Wang, 2023), and Sun Simiao-7B(Xin Yan, 2023), as baselines. Moreover, we incorporate two state-of-the-art LLMs i.e., GPT-4o(Hurst et al., 2024) (one of the most advanced LLMs) and Qwen-Max(Team, 2024) (a leading Chinese-language LLM) for comparison. Various evaluation metrics are designed to ensure a comprehensive and fair assessment. Different tasks employed task-specific evaluation criteria, enabling a multidimensional analysis of the performance of JF.
@@ 4.2 Assessment of Syndrome Differentiation Accuracy
We have preprocessed over 63,000 real diagnostic data entries to select over 43,000 high-quality data entries for LLM fine-tuning. Based on this data, the Qwen2.5-7B-Instruct model and the Roberta model are taken respectively as the foundation model of JF for training its syndrome differentiation ability. Here, syndrome differentiation can also be considered as a classification task, thus we train it via these two different types of models. For the evaluation of TCM syndrome differentiation accuracy, we select 8,699 real-world TCM cases as the test dataset, covering 170 different syndrome types. Considering the imbalance in sample sizes across different categories, we use weighted evaluation metrics to assess the model’s performance in various categories. The metrics include weighted precision: $P_{w}$, weighted recall: $R_{w}$, and weighted F1: $F1_{w}$, which are as follows:
$P_{w}=\frac{\sum_{i=1}^{n}w_{i}\cdot TP_{i}}{\sum_{i=1}^{n}w_{i}\cdot(TP_{i}+FP_{i})},$ (13)
$R_{w}=\frac{\sum_{i=1}^{n}w_{i}\cdot TP_{i}}{\sum_{i=1}^{n}w_{i}\cdot(TP_{i}+FN_{i})},$ (14)
$F1_{w}=\frac{\sum_{i=1}^{n}w_{i}\cdot TP_{i}}{\sum_{i=1}^{n}w_{i}\cdot(TP_{i}+0.5\cdot(FP_{i}+FN_{i}))}.$ (15)
Here, $n$ is the number of TCM syndrome categories. $w_{i}$ represents the weight of the $i$-th category, which is calculated as the number of samples in category $i$ divided by the total number of samples. $TP_{i}$ is the number of true positives for category $i$, $FP_{i}$ is the number of false positives for category $i$, and $FN_{i}$ is the number of false negatives for category $i$.
Model $P_{w}$ $R_{w}$ $F1_{w}$ JingFang-RoBERTa 0.8185 0.8230 0.8186 JingFang-Qwen-2.5-7B 0.8015 0.8145 0.8032 GPT-4o 0.4328 0.2138 0.2474 Qwen-max 0.4359 0.1811 0.2249 Sun Simiao 0.3113 0.1497 0.1415 Zhong Jing 0.0715 0.0598 0.0308 Ming Yi 0.1380 0.0467 0.0217 Shen Nong 0.0713 0.0945 0.0401
Table 1: Comparison of TCM syndrome differentiation capabilities among different models.
Proactivity Accuracy Practicality Overall Effectiveness Total Score
Jing Fang 8.45 8.19 8.44 8.42 8.38
Sun Simiao 6.56 5.50 5.41 5.77 5.74
Shen Nong 6.56 5.50 5.47 5.95 5.92
Ming Yi 4.44 5.19 4.88 5.14 4.85
Zhong Jing 4.31 4.50 4.31 3.80 4.20
Table 2: Comparison of multi-round consultations capabilities across different models.
As shown in Table 1, JF surpasses both the baseline models and general models across all evaluation metrics in the TCM syndrome differentiation task. Specifically, the JingFang-RoBERTa model achieved a precision of 0.8185, a recall of 0.8230, and a F1 score of 0.8186. Notably, the general models GPT-4o and Qwen-Max, due to their significantly larger number of parameters compared to the baseline models, outperformed other baseline models in certain metrics. However, their performance still does not reach the level of ours. These results demonstrate that our models possess the expert-level capability of syndrome differentiation. Here, we need to emphasize that the choice of foundation models is not a factor affecting its capabilities in syndrome differentiation, while the training method and data do. For more information, see the Ablation Experiment and appendix.
@@ 4.3 Medical Consultation Evaluation
For the evaluation of medical consultation ability, we randomly selected 100 cases from real TCM medical cases. These cases involve a variety of clinical conditions, which comprehensively reflect the diversity and complexity of TCM consultation. Based on these cases, an LLM is used to act as the patient and interact with JF and each baseline model to complete multi-round consultation, collecting the patient’s medical information.
To professionally assess the capacity of medical consultation in multi-round consultation, multiple TCM experts are invited to score the models’ performance from 0-10 points respectively based on four dimensions: Proactiveness, Accuracy, Practicality, and overall effectiveness, with a total score of 40 points. (For the specific evaluation criteria for each dimension, refer to Sec.A.) The evaluation in each dimension carefully considers the model’s key abilities, such as proactiveness in gathering patient information and the model’s responsiveness to the patient’s actual needs in TCM consultations.
As shown in Table 2, our MDCCTM significantly advances LLM’s medical consultation capacity, and JF outperforms all baseline models. The scores for Proactiveness, accuracy, practicality, and overall effectiveness all exceeded 8, notably surpassing other models. These results indicate the MDCCTM, which is a CoT-driven multi-agent collaborative mechanism, exhibits superior interaction and information processing capabilities during multi-round consultations. According to the performance, MDCCTM provides a foundation for LLM of more accurate clinical diagnosis and highlights JF’s potential for real-world applications.
@@ 4.4 Ablation Experiment
4.4.1 Evaluation of TCM General Agent in Medical Consultation
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MDCCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation. This experiment selects 100 representative TCM cases from real medical data, ensuring typicality and diversity. To simulate multi-round interactions, we employ an LLM-based patient agent, which dynamically generates consultation responses from a patient’s perspective, effectively mimicking real-world consultation scenarios. To evaluate performance, we conducted a comparative analysis, assessing consultation results against the original case information. The evaluation focuses on the comprehensiveness (coverage of multi-dimensional health information) and the pertinence (adaptation to case-specific features). To ensure objectivity, GPT-4o is used as the evaluation tool to automatically compare consultation outputs from both configurations and score them based on these criteria.
@@ 4.4.2 Applicability Assessment of the TCM Syndrome Differentiation Dataset
This ablation experiment is designed to demonstrate the effectiveness and applicability of our proposed method and constructed data for TCM syndrome differentiation and explore the influence of different foundation models (or the model with different sizes of parameters) on syndrome differentiation. We chose multiple representative open-source LLMs to test, which include the Qwen-2.5 series (Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, Qwen-2.5-14B-Instruct)(Yang et al., 2024a), the DeepSeek series (DeepSeek-7B-Chat)(DeepSeek-AI, 2024), and the Llama series (Llama3.1-8B-Instruct)(Dubey et al., 2024).
As shown in Figure 5, all models in the test exhibit advanced performance in TCM syndrome differentiation compared to their original version before being trained via our method and data. According to the results, we observe that the choice of foundation model is not the key factor for improving the syndrome differentiation ability, while the method and data do. Moreover, compared to the performance of GPT-4o (0.4328), the results highlight the potential applicability and value of the proposed TCM syndrome differentiation method and dataset for future research. For a detailed description of the ablation experiment, refer to Sec.D.
@@ 5 Conclusion and Discussion
Moreover, continuing to deeply explore the potential application of the multi-agent collaborative mechanism and enhancing the innovative application of LLMs in different fields will be interesting. At the same time, in the field of TCM LLM, developing advanced multi-modal TCM large models to promote more efficient application of TCM, will also be very meaningful.
Impact Statement
This paper presents work that advances the application field of LLMs, particularly in the TCM domain. The proposed model can greatly improve the efficiency of doctors in medical diagnosis and treatment. However, in addition to existing ethical questions for this area (e.g. inaccurate content, bias, toxicity), our model cannot completely replace the role of doctors in the real clinical practice of TCM. Due to individual differences in patients and the complexity of real medical scenarios, specific diagnoses and treatments require careful adherence to the advice of real doctors.
Appendix A TCM Consultation
To enable efficient consultation with the patient, we propose an ”Intelligent Dynamic Consultation Information Collection and Termination” Algorithm. This algorithm automatically collects key points from the consultation CoT through a limited number of consultation rounds and automatically terminates the dialogue based on this consultation CoT when specific conditions are met. Based on TCM consultation knowledge, the key points are categorized into four types: symptom information, medical history information, lifestyle information, and other information, which helps to systematically extract crucial details. The algorithm’s inputs include the maximum consultation rounds, patient inputs, and the consultation CoT, and through these inputs, it generates a set of key points and drives the conversation forward. Finally, the algorithm outputs the collected set of medical information points.
@@ Appendix A TCM Consultation
Appendix B TCM Syndrome Differentiation
B.1 Dataset Construction for Syndrome Differentiation
@@
During the construction of the syndrome differentiation dataset, we observed that real-world TCM case and diagnostic data often contain a substantial amount of information irrelevant to the patient’s condition, which cannot be easily removed through regularization techniques. For instance, phrases such as ”The patient visited our hospital for further integrated traditional Chinese and Western medicine treatment” or ”The patient was advised to undergo cranial CT and blood tests but refused” are not only extraneous to the core condition but also hinder the model’s ability to focus on key information about the illness. Moreover, the raw TCM case and diagnostic data frequently exhibit a high degree of redundancy in describing the patient’s condition and physical signs, resulting in a low density of valid information, which adversely impacts fine-tuning performance and the generalization ability of LLMs.
To address these issues, we proposed a universal syndrome differentiation data cleaning method based on the TCM framework of the ”Ten Questions Song.” This strategy aims to simulate the professional judgment of TCM experts using LLMs, accurately screening and reconstructing raw TCM case and diagnostic data to construct a high-quality syndrome differentiation dataset. Specifically, this approach strictly retains core information closely related to the patient’s condition, optimizes the medical value of the data, and organizes the content systematically according to the ”Ten Questions Song” framework to ensure the dataset includes key elements required for TCM syndrome differentiation. Through this strategy, we effectively eliminate redundant information, increase the structuralization of the data, and better adapt it to the application requirements of LLMs in TCM syndrome differentiation tasks.
As shown in Figure 6, the data-cleaning process enforces strict adherence to the ”Ten Questions Song” framework by applying explicit data extraction rules, retaining only information directly related to the patient’s condition. A comparison of the raw TCM case and diagnostic data with the cleaned dataset (Preconditioned TCM Data) reveals that the cleaned records are more focused on core symptoms, physical signs, and the evolution of the condition. Unrelated information is systematically removed, resulting in texts that are both precise and medically valuable. The core optimization points of this method include the following:
@@
    Strengthening the comprehensiveness and pertinence of syndrome differentiation dataset: Comprehensiveness is reflected in the inclusion of all key consultation dimensions from the TCM ”Ten Questions Song,” such as cold/heat, sweating, head/body, bowel movements, diet, sleep, medical history, and family history, ensuring the completeness of diagnostic criteria. Pertinence is achieved by prioritizing the retention of information related to core elements of syndrome differentiation (e.g., chief complaints, symptom changes, disease progression, and physical signs) while removing background or secondary information. This makes the dataset more aligned with the logic of TCM syndrome differentiation, thereby improving the accuracy of LLMs in clinical reasoning.
@@
B.2 Model Training for Syndrome Differentiation
@@
Based on the syndrome differentiation data-cleaning method proposed in this study, a total of over 43,000 high-quality syndrome differentiation data entries were obtained. These entries were split into training and testing datasets at a ratio of 8:2. Subsequently, the Qwen2.5-7B-Instruct model was fine-tuned using the LoRA method. The form of the training dataset is shown in Figure 7. In the experiments, we employed four A6000 GPUs (each with 48GB of memory). After multiple rounds of experimental tuning, we determined the optimal training parameters to ensure the stable convergence of the loss function. The specific training parameters included: ”lora rank”: 64, ”lora alpha”: 16, ”learning rate”: ”5e-4”, and ”batch size”: 16.
@@
Appendix C TCM Treatment Recommendation
Syndrome differentiation-based treatment is one of the core concepts in TCM. The fundamental idea is to identify the etiology and pathogenesis of a patient’s condition and then adopt a personalized treatment strategy. Different diseases may present similar syndrome differentiation; for instance, the syndrome of wind-cold attacking the lungs may manifest not only in the common cold but also in acute bronchitis. Therefore, experienced TCM practitioners often rely on the results of syndrome differentiation, integrating the patient’s medical history and constitution, to provide individualized treatment recommendations. This study employs a Dual-Stage approach to recall relevant TCM treatment knowledge based on the TCM formula dataset we have constructed.
In the first stage, the system screens candidate prescription data related to the patient’s syndrome type from the TCM prescription database based on the differentiation outcomes. This screening process utilizes the patient’s syndrome differentiation information, including syndrome type, etiology, and disease location, to narrow down the candidate prescriptions to those highly relevant to the patient’s condition, thereby ensuring the pertinence and effectiveness of subsequent treatment recommendations. In the second stage, the system employs an embedding method to vectorize the patient’s detailed medical history information and the clinical manifestations of the candidate prescriptions, subsequently calculating the similarity between them. This method allows for the assessment of how well the candidate prescriptions match the patient’s condition. Finally, based on similarity rankings, the top three cases most similar to the patient’s medical history information are selected (TOP-3). These ”top three most similar cases” refer to cases that closely resemble the current patient’s condition across multiple dimensions, including syndrome differentiation information, medical history, and symptom presentation. The detailed methodology is presented in Algorithm 4.
Algorithm 4 MDCCTM for TCM Treatment Recommendation

### v2

@@ 3 Experiments
In this section, we demonstrate the performance of JF on two core aspects: the accuracy of TCM syndrome differentiation and the comprehensiveness and effectiveness of clinical consultation, via a series of systematic experiments.
3.1 Datasets and Baseline Model
The dataset comes from three parts: the open-source TCM-SD [22] syndrome differentiation dataset, 48 modern TCM textbooks covering internal medicine, surgery, gynecology, and pediatrics, 700 ancient TCM texts. We manually cleaned the data and engaged three PhD students from different specialties at Beijing University of Chinese Medicine to proofread the data. We have constructed a high-quality TCM knowledge database for accurate treatment recommendation which consists of 80,000 prescription entries, including 1,000 classic formulas. Each entry contains key attributes such as disease type, syndrome type, clinical manifestations, syndrome mechanism, treatment methods, representative formulas, and commonly used medicines. We also have processed over 63,000 records of syndrome differentiation in raw data and selected 43,000 high-quality samples with detailed case descriptions and corresponding diagnoses for model training on syndrome differentiation.
To comprehensively evaluate the performance of JF, we selected several representative open-source TCM models, i.e., Zhongjing-7B [31], HuaTuo-32B [5], ShenNong-7B [28], and SunSimiao-7B [29], as baselines. Moreover, we incorporate four state-of-the-art (SOTA) LLMs i.e., Glaude-3.7 [2], DeepSeek v3 [10], Gemini-1.5-pro[23] and GPT-4o [18] for comparison. Various evaluation metrics are designed to ensure a comprehensive and reliable assessment. Different tasks adopts task-specific evaluation criteria, enabling a multidimensional analysis for the performance of JF.
@@ 3.2 Assessment of Syndrome Differentiation Accuracy
Based on the curated TCM syndrome differentiation data, the Qwen2.5-7B-Instruct model and the Roberta model are taken respectively as the foundation model of JF for training its syndrome differentiation ability. Here, syndrome differentiation can also be considered as a classification task, thus we train it via these two different types of models. The training was conducted on four A6000 GPUs, each with 48GB of VRAM. For the evaluation of syndrome differentiation accuracy, we randomly select 8,699 TCM cases as the test dataset, covering 170 different syndrome types. As the imbalance in sample sizes across different categories, we use weighted evaluation metrics to assess the model’s performance for the sake of fair and accurate evaluation. The metrics include weighted precision: $P_{w}$, weighted recall: $R_{w}$, and weighted F1: $F1_{w}$.
  Model $P_{w}$ $R_{w}$ $F1_{w}$     Model $P_{w}$ $R_{w}$ $F1_{w}$
JF-RoBERTa 0.8185 0.8230 0.8186     Glaude-3.7 0.2731 0.1206 0.1453
JF-Qwen-2.5-7B 0.8015 0.8145 0.8032     HuaTuo 0.3654 0.1735 0.1515
DeepSeek v3 0.4700 0.3032 0.3221     SunSimiao 0.3113 0.1497 0.1415
Gemini-1.5-pro 0.4532 0.1992 0.2209     ShenNong 0.0713 0.0945 0.0401
GPT-4o 0.4328 0.2138 0.2474     ZhongJing 0.0715 0.0598 0.0308
Table 1: Comparison of TCM syndrome differentiation performance across models.
As shown in Table 1, JF surpasses both the baseline models and the SOTA general models across all evaluation metrics in the TCM syndrome differentiation task. Specifically, JF model improves at least 74.1% in accuracy compared to the SOTA general model (DeepSeek-V3), and improves at least 124% in accuracy compared to the TCM model (HuaTuo-32B). Notably, the general models Glaude-3.7, DeepSeek v3, Gemini-1.5-pro and GPT-4o, due to their significantly larger number of parameters compared to the baseline models, outperformed other baseline models in certain metrics. However, their performance still does not reach the level of JF. These results demonstrate that our models possess the expert-level capability of syndrome differentiation. Here, we need to emphasize that the choice of foundation models is not a factor affecting its capabilities in syndrome differentiation, while the training method and data do. For more information, see the section 3.4.2
@@ 3.3 Medical Consultation Evaluation
For the evaluation of clinical consultation ability, we randomly selected 50 TCM medical cases from the database. These cases involve a variety of clinical conditions, which comprehensively reflect the diversity and complexity of TCM consultation. An LLM is used to act as the patient based on these cases, interacting with JF, baseline models and SOTA general models to accomplish the multi-round consultation and collect the patient’s medical information.
(a) Evaluation results for Proactivity
(b) Evaluation results for Accuracy
(c) Evaluation results for Practicality
(d) Evaluation results for Overall Effectiveness
Figure 3: Comparison of multi-round consultation capabilities across different models.
To comprehensively evaluate the multi-round medical consultation capabilities, 10 TCM experts were invited to assess the performance of each model across four key dimensions: Proactiveness, Accuracy, Practicality, and Overall Effectiveness. Pairwise comparisons are conducted between the proposed JF, each baseline and SOTA general model. Based on performance, each comparison was categorized as a win, loss, or tie (For detailed evaluation criteria of each dimension, refering to section B). The assessment in each dimension focused on the model’s performance in critical consultation skills, comprehensiveness and pertinence.
As shown in Figure 3, the MACCTM significantly improves the model’s overall performance on clinical consultation. Compared with various baseline and general models, the JF consistently outperforms all comparison models across the four dimensions of proactivity, accuracy, practicality, and overall effectiveness. These results indicate the MACCTM, which is a CoT-driven multi-agent collaborative mechanism, exhibits superior interaction and information processing capabilities during multi-round consultations. Based on performance, MACCTM provides a solid foundation for model to carry out accurate clinical diagnosis and highlights its potentiality in real-world applications.
@@ 3.4 Ablation Experiment
3.4.1 Evaluation of TCM General Agent MACCTM
To evaluate the impact of the combination of TCM Specialist Agent and TCM General Agent in MACCTM on multi-round medical consultation and to analyze the key role and contributions of the TCM General Agent, this study compares two configurations: JF without TCM General Agent and JF with TCM General Agent, systematically analyzing their differences in performance in multi-round consultation. This experiment selects 100 representative TCM cases from medical data, ensuring typicality and diversity. To simulate multi-round interactions, we employ an LLM-based patient agent, which dynamically generates consultation responses from a patient’s perspective, effectively mimicking real-world consultation scenarios. To evaluate performance, we conducted a comparative analysis, assessing consultation results against the original case information. The evaluation focuses on the comprehensiveness (coverage of multi-dimensional health information) and the pertinence (adaptation to case-specific features). To ensure objectivity, GPT-4o is used as the evaluation tool to automatically compare consultation outputs from both configurations and score them based on these criteria.
@@ 3.4.2 Applicability Assessment of the TCM Syndrome Differentiation Dataset
This ablation experiment is designed to demonstrate the effectiveness and applicability of our proposed method and constructed data for TCM syndrome differentiation and explore the influence of different foundation models (or the model with different size of parameters) on syndrome differentiation. We chose several representative open-source LLMs to test, which include the Qwen-2.5 series (Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Instruct, Qwen-2.5-14B-Instruct) [30], the DeepSeek series (DeepSeek-7B-Chat) [9], and the Llama series (Llama3.1-8B-Instruct) [12], and the GLM series (GLM-4-9B Chat) [13], and the Gemma series (Gemma-3-12b-it) [24].
As shown in Figure 5, all models in the experiment exhibit advanced performance in TCM syndrome differentiation compared to their original version before being trained via our method and data. According to the results, we observe that the choice of foundation model is not the key factor for improving the syndrome differentiation ability, while the method and data do. Moreover, compared to the performance of GPT-4o (0.4328 in precison), the results highlight the potential applicability and value of the proposed TCM syndrome differentiation method and dataset for future research. For a detailed description of the ablation experiment, refer to Sec.E.
@@ 4 Conclusion and Discussion
Appendix
Appendix A Framework of JingFang
In this section, we introduce the framework of the proposed JingFang (referred to as the JF framework). As shown in Figure 1, following the authentic process of TCM clinical consultation, diagnosis, and treatment, JF framework is designed to consist of three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment Reconmmendation. It follows the systematically designed MACCTM and DSRS to fulfill the professional and accurate capabilities of consultation and syndrome differentiation-based treatment.
TCM Consultation Module mainly involves the following steps:
① Expert Team Construction: The Medical Record Agent summarizes the patient’s symptoms and basic information to generate the chief complaint. Based on the chief complaint, the relevant TCM Specialist Agents and a TCM General Agent are gathered to form the TCM expert team for the diagnosis.
TCM Treatment Recommendation Module provides professional and targeted treatment recommendations for patients. The Treatment Agent is generated through a DSRS to provide personalized treatment suggestions based on the patient’s syndrome type and the patient’s condition.
Appendix B TCM Consultation
To enable efficient consultation with the patient, we propose an "Intelligent Dynamic Consultation Information Collection and Termination" Algorithm. This algorithm automatically collects key points from the consultation CoT through a limited number of consultation rounds and automatically terminates the dialogue based on this consultation CoT when specific conditions are met. Based on TCM consultation knowledge, the key points are categorized into four types: symptom information, medical history information, lifestyle information, and other information, which helps to systematically extract crucial details. The algorithm’s inputs include the maximum consultation rounds, patient inputs, and the consultation CoT, and through these inputs, it generates a set of key points and drives the conversation forward. Finally, the algorithm outputs the collected set of medical information points.
@@ Appendix B TCM Consultation
Appendix C TCM Syndrome Differentiation
C.1 Dataset Construction for Syndrome Differentiation
@@
During the construction of the syndrome differentiation dataset, we observed that TCM cases in the dataset often contain a substantial amount of information irrelevant to the patient’s condition, which cannot be easily removed through regularization techniques. For instance, phrases such as "The patient visited our hospital for further integrated traditional Chinese and Western medicine treatment" or "The patient was advised to undergo cranial CT and blood tests but refused" are not only extraneous to the core condition but also hinder the model’s ability to focus on key information about the illness. Moreover, the raw TCM case and diagnostic data frequently exhibit a high degree of redundancy in describing the patient’s condition and physical signs, resulting in a low density of valid information, which adversely impacts fine-tuning performance and the generalization ability of LLMs.
To address these issues, we proposed a universal syndrome differentiation data cleaning method based on the TCM framework of the "Ten Questions Song." This strategy aims to simulate the professional judgment of TCM experts using LLMs, accurately screening and reconstructing raw TCM case and diagnostic data to construct a high-quality syndrome differentiation dataset. Specifically, this approach strictly retains core information closely related to the patient’s condition, optimizes the medical value of the data, and organizes the content systematically according to the "Ten Questions Song" framework to ensure the dataset includes key elements required for TCM syndrome differentiation. Through this strategy, we effectively eliminate redundant information, increase the structuralization of the data, and better adapt it to the application requirements of LLMs in TCM syndrome differentiation tasks.
As shown in Figure 6, the data-cleaning process enforces strict adherence to the "Ten Questions Song" framework by applying explicit data extraction rules, retaining only information directly related to the patient’s condition. A comparison of the raw TCM case and diagnostic data with the cleaned dataset (Preconditioned TCM Data) reveals that the cleaned records are more focused on core symptoms, physical signs, and the evolution of the condition. Unrelated information is systematically removed, resulting in texts that are both precise and medically valuable. The core optimization points of this method include the following:
@@
    Strengthening the comprehensiveness and pertinence of syndrome differentiation dataset: Comprehensiveness is reflected in the inclusion of all key consultation dimensions from the TCM "Ten Questions Song," such as cold/heat, sweating, head/body, bowel movements, diet, sleep, medical history, and family history, ensuring the completeness of diagnostic criteria. Pertinence is achieved by prioritizing the retention of information related to core elements of syndrome differentiation (e.g., chief complaints, symptom changes, disease progression, and physical signs) while removing background or secondary information. This makes the dataset more aligned with the logic of TCM syndrome differentiation, thereby improving the accuracy of LLMs in clinical reasoning.
@@
C.2 Model Training for Syndrome Differentiation
@@
Based on the syndrome differentiation data-cleaning method proposed in this study, a total of over 43,000 high-quality syndrome differentiation data entries were obtained. These entries were split into training and testing datasets at a ratio of 8:2. Subsequently, the Qwen2.5-7B-Instruct model was fine-tuned using the LoRA method. The form of the training dataset is shown in Figure 7. In the experiments, we employed four A6000 GPUs (each with 48GB of memory). After multiple rounds of experimental tuning, we determined the optimal training parameters to ensure the stable convergence of the loss function. The specific training parameters included: "lora rank": 64, "lora alpha": 16, "learning rate": "5e-4", and "batch size": 16.
@@
Appendix D TCM Treatment Recommendation
Syndrome differentiation-based treatment is one of the core concepts in TCM. The fundamental idea is to identify the etiology and pathogenesis of a patient’s condition and then adopt a personalized treatment strategy. Different diseases may present similar syndrome differentiation; for instance, the syndrome of wind-cold attacking the lungs may manifest not only in the common cold but also in acute bronchitis. Therefore, experienced TCM practitioners often rely on the results of syndrome differentiation, integrating the patient’s medical history and constitution, to provide individualized treatment recommendations. This study employs a Dual-Stage approach to recall relevant TCM treatment knowledge based on the TCM formula dataset we have constructed. An example from the TCM formula database is shown in the Figure 9.
Figure 9: Example from the TCM Formula Database
In the first stage, the system screens candidate prescription data related to the patient’s syndrome type from the TCM prescription database based on the differentiation outcomes. This screening process utilizes the patient’s syndrome differentiation information, including syndrome type, etiology, and disease location, to narrow down the candidate prescriptions to those highly relevant to the patient’s condition, thereby ensuring the pertinence and effectiveness of subsequent treatment recommendations. In the second stage, the system employs an embedding method to vectorize the patient’s detailed medical history information and the clinical manifestations of the candidate prescriptions, subsequently calculating the similarity between them. This method allows for the assessment of how well the candidate prescriptions match the patient’s condition. Finally, based on similarity rankings, the top three cases most similar to the patient’s medical history information are selected (TOP-3). These "top three most similar cases" refer to cases that closely resemble the current patient’s condition across multiple dimensions, including syndrome differentiation information, medical history, and symptom presentation. The detailed methodology is presented in Algorithm 3. The "sparse" and "dense" embeddings in Algorithm 3 are two complementary vector representations in our Dual-Stage Retrieval Scheme: sparse embeddings (emb_sparse) are high-dimensional vectors with mostly zero elements, capturing exact lexical matches and keyword presence in medical records, excelling at precise identification of medical terminology and symptom descriptors; dense embeddings (emb_dense) are low-dimensional vectors with semantics distributed across all dimensions, capturing semantic relationships and contextual understanding beyond keyword matches, helping identify conceptually relevant prescriptions.
Algorithm 3 DSRS for TCM Treatment Recommendation